OpenStudy (anonymous):

which equations correctly represent a line that has a slope of -2/3 and passes through the points (–2, 8) and (1, 6)? A. y – 2 = -2/3(x + 8) B. y + 2 = -2/3(x – 8) C. y – 6 = -2/3(x – 1) D. y – 1 = -2/3(x – 6)

OpenStudy (anonymous):

use this formula \ (y-y1) = m (x-x1) this above formula is called point slope form in the question slope is given m=-2/3 now u can take any one point from the above question's two points{ (-2,8) , (1,6) } so i take the ( 1,6 ) now just put it in the above formula

OpenStudy (anonymous):

@rio96 is it A ?

OpenStudy (anonymous):

(y-y1)=m(x-x1) m=-2/3 i choose randomly this point (1,6) okay? now i m going to put it into the above formula y-6 =-2/3 (x-1) i solved this question the correct answer is C